On April 1, 2013, Guyton Sails accepted a $12,000 advance payment for maintaining a customer’s fleet...

  1. On April 1, 2013, Guyton Sails accepted a $12,000 advance payment for maintaining a customer’s fleet of sail boats. The contract is for a twelve month period. What is the remaining liability that would appear on the company’s balance sheet at the end of May?

    $10,000

    $1,200

    $12,000

    $10,800
Answer #1

Answer:

Correct answer is:

$10,000

Explanation:

Earned revenue each month = 12000 / 12 = $1,000

Earned till May end (for April and May) = 1000 * 2 = $2,000

Remaining liability at the end of May = 12000 - 2000 = $10,000

Hence option A is correct and other options B, C and D are incorrect.